# Volatility Spillovers and Interconnectedness in OPEC Oil Markets: A Network-Based log-ARCH Approach

### v1

@@
Volatility Spillovers and Interconnectedness in OPEC Oil Markets: A Network-Based log-ARCH Approach
@@
   Kahina Mehidi Department of Economics, University of Bejaia, Algeria. Email: kahina.mehidi@univ-bejaia.dz    Khelifa Mazouz Department of Economics, University of Bejaia, Algeria. Email: khelifa.mazouz@univ-bejaia.dz;
Cardiff Business School, Cardiff University, United Kingdom. Email: MazouzK@cardiff.ac.uk
   Philipp Otto School of Mathematics and Statistics, University of Glasgow, United Kingdom. Email: Philipp.Otto@glasgow.ac.uk; ORCID: $0$-0002-9796-6682
(July 20, 2025)
This paper examines several network-based volatility models for oil prices, capturing spillovers among OPEC oil-exporting countries by embedding novel network structures into ARCH-type models. We apply a network-based log-ARCH framework that incorporates weight matrices derived from time-series clustering and model-implied distances into the conditional variance equation. These weight matrices are constructed from return data and standard multivariate GARCH model outputs (CCC, DCC, and GO-GARCH), enabling a comparative analysis of volatility transmission across specifications. Through a rolling-window forecast evaluation, the network-based models demonstrate competitive forecasting performance relative to traditional specifications and uncover intricate spillover effects. These results provide a deeper understanding of the interconnectedness within the OPEC network, with important implications for financial risk assessment, market integration, and coordinated policy among oil-producing economies.
@@ 1 Introduction
The Organization of the Petroleum Exporting Countries (OPEC) plays a central role in stabilizing global oil markets by coordinating production and pricing among its members. Acting as a cartel, OPEC restricts oil supply to support prices (Kilian, 2009; Al Rousan et al., 2018). Despite this coordination, significant differences persist across member countries regarding production behavior and pricing dynamics. Understanding how volatility transmits across these economies remains an important, unresolved question.
Recent studies have increasingly adopted network-based methodologies to explore volatility spillover and interdependencies. These include dynamic production networks (Al Rousan et al., 2018), time-varying spillover indices (Ji et al., 2018), multilayer spillover structures (Dai et al., 2023), and global energy networks (Uddin et al., 2023). Many of these rely on the connectedness framework of Diebold and Yilmaz (2009, 2012, 2014), which measures directional spillovers using forecast error variance decompositions. However, these approaches often depend on external explanatory variables or impose structure through rolling-window VARs, and do not directly model the volatility process or fully account for the network-based structure of conditional heteroskedasticity.
This paper addresses this gap by asking: Can volatility spillovers among OPEC countries be better captured by embedding network structures directly into volatility models? Our motivation stems from the need for models that reflect both the time-varying nature of volatility and the structural interdependence among oil-exporting countries. In particular, the OPEC framework presents a unique case in which policy coordination coexists with heterogeneity in market responses. This calls for a modeling strategy capable of uncovering latent transmission channels that are both temporal and cross-sectional.
To address this, we develop a network-based volatility modeling framework. Building on the spatiotemporal ARCH modeling literature, such as Otto et al. (2024a), we introduce a methodology that embeds time-series clustering-derived network structures into the volatility specification. This allows us to model volatility transmission not only over time but also across a network of interdependent economies. Our approach does not rely on external variables; instead, it infers spillover dynamics directly from the historical log return data.
We focus on six OPEC countries, selected based on the completeness of their monthly historical log return data from the official OPEC database. We construct several dissimilarity-based weight matrices using raw data Euclidean distances, correlations, and model-implied metrics from CCC-GARCH (Bollerslev, 1990), DCC-GARCH (Engle, 2002), GO-GARCH (Van der Weide, 2002), and the financial networks considered in the network log-ARCH models of Mattera and Otto (2024). These weight matrices define the network structures used in the log-ARCH specification. The model is implemented within a rolling-window forecasting scheme to evaluate out-of-sample performance.
Our empirical results show that network-based log-ARCH models incorporating CCC, DCC, and GO-GARCH-derived weight matrices demonstrate strong forecasting performance relative to standard alternatives and substantially reduce the computational complexity associated with the standard multivariate GARCH models. These models yield lower Root Mean Squared Forecast Errors (RMSFE) and Mean Absolute Forecast Errors (MAFE) than most competing specifications. More importantly, they uncover richer and more detailed spillover patterns across the OPEC network. These patterns often align with known geopolitical and institutional ties, even though no such data are explicitly included in the model.
In addition to improving forecast accuracy, the proposed framework enhances interpretability. This reveals the latent transmission channels of volatility that reflect the evolving relationships between member states. For investors, this provides a tool for improved volatility forecasting and dynamic risk management. For policymakers, the model offers early warning signals of breakdowns in coordination or emerging systemic risks. Since the method relies only on return data, it enables real-time assessment without the need for external indicators.
@@
The remainder of this paper is organized as follows. Section 2 reviews the relevant literature on oil price volatility, forecasting, and network-based modeling. Section 3 describes the dataset and presents the key descriptive statistics. Section 4 provides an overview of univariate and multivariate volatility models. Section 5 outlines the methodological framework for the network-based log-ARCH models and the construction of the network weight matrices. Section 6 presents the empirical findings, including the estimation results, network structures highlighting interconnectedness and spillover effects, and the comparative forecasting performance of the proposed models.
@@ 2 Literature Review
Since its establishment in Baghdad in 1960, OPEC has aimed to coordinate and harmonize the petroleum policies of its member countries to ensure stable oil prices and adequate returns on investment (OPEC Statute, ). Over the decades, the organization has been shaped by a series of major historical and geopolitical events (OPEC Brief History, ), many of which have had far-reaching effects on the global economy (Jo, 2014). These events have exposed the persistent susceptibility of oil-exporting countries to external shocks, highlighting the complexity of managing coordinated policies in the face of heterogeneous economic responses (Liu et al., 2019). Geopolitical risks, particularly in oil-producing regions such as the Middle East and North Africa, have consistently driven oil price volatility (Mignon and Saadaoui, 2024). For instance, regional tensions and conflicts, along with global crises—such as the Chinese economic slowdown (Hamilton, 2014)—have led to sharp fluctuations in oil prices. More recently, the COVID-19 pandemic amplified this volatility by triggering both supply- and demand-side shocks (Bourghelle et al., 2021; Le et al., 2023). Such compounded disruptions underscore the importance of understanding how volatility transmits across oil-exporting economies, particularly within policy-coordinated networks such as OPEC, where shocks may propagate unevenly across members.
@@
Oil price volatility remains a central concern in economic and financial forecasting, given its influence on key macroeconomic indicators, such as investment, durable goods consumption, and aggregate output. Uncertainty generated by oil price fluctuations can lead to reduced economic activity, particularly in resource-rich exporting nations (Finn, 2000; Kilian, 2008; Elder and Serletis, 2010; Vo and Le, 2021). These countries, as emphasized by Cappelli et al. (2023), are disproportionately affected due to their strong fiscal reliance on oil revenues. Despite decades of research, accurately forecasting oil price movements remains a formidable task due to complex interactions among market agents (Baumeister and Kilian, 2016a).
@@
Conventional econometric models often fail to fully capture the complexity of oil price volatility, particularly its dynamic and interconnected nature within coordinated economic blocs. In response, Engle (1982) and Bollerslev (1986) developed the ARCH and GARCH models, which model volatility clustering in financial time series. These nonlinear frameworks remain foundational in financial econometrics, especially through multivariate extensions such as CCC-GARCH (Bollerslev, 1990), DCC-GARCH (Engle, 2002), and GO-GARCH (Van der Weide, 2002), which accommodate dynamic correlations in the return series. More recently, attention has shifted to models that incorporate spatial and network-based dependencies to account for cross-country volatility transmission. Spatiotemporal ARCH (spARCH) models incorporate contemporaneous spatial and temporal dependence to better capture volatility interlinkages (Otto et al., 2018), while spatial and spatiotemporal log-ARCH models (Sato and Matsuda, 2021; Otto et al., 2024a) and network log-ARCH models (Mattera and Otto, 2024) extend this framework to account for multiplicative effects. A comprehensive synthesis of these advanced approaches is presented in Otto et al. (2024b).
Empirical studies have shown strong interdependencies between oil prices and financial markets (Awartani and Maghyereh, 2013; Guhathakurta et al., 2020; Guan et al., 2024). For instance, Nguyen et al. (2020) document the short-run global repercussions of oil price shocks, while Creti et al. (2014) observe heightened oil–stock market co-movements in oil-exporting countries. However, existing studies often rely on models that assume static or exogenous relationships, limiting their ability to reflect the endogenous and evolving nature of the interdependencies among oil-exporting countries. In response, recent advances—such as time-series clustering methods (Maharaj et al., 2019) embedded into Dynamic Spatiotemporal ARCH models (Otto et al., 2024a)—offer promising tools for inferring network structures directly from return data. This opens new avenues for modeling volatility transmission in a way that reflects both temporal dynamics and cross-sectional spillovers, which is particularly relevant for OPEC economies, where coordinated production masks heterogeneous national responses. By integrating clustering-derived networks with advanced volatility models, our approach captures the latent and evolving interdependencies that shape volatility spillovers in oil-exporting nations. These methods move beyond traditional frameworks by eliminating the need for external variables, allowing for real-time analysis rooted in the structure of the return data itself. This offers both improved forecast accuracy and deeper insights into the underlying transmission channels relevant for policy coordination and risk management.
@@ 3 Data, Descriptive Statistics and Preliminary Analyses
The empirical analysis in this study is based on monthly crude oil price data for six key OPEC member countries: Algeria (Saharan Blend), Iran (Heavy), Libya (Es Sider), Nigeria (Bonny Light), Saudi Arabia (Arab Light), and the United Arab Emirates (Murban). These countries were selected based on the availability of a complete and consistent historical time series, which offered the same number of monthly observations across the entire sample. Crude oil prices were retrieved from the official OPEC database and subsequently converted into monthly logarithmic returns. The dataset spans the period from January 1983 to December 2022.
@@
Figure 1 presents the time series of the logarithmic returns and their squared values, capturing the volatility dynamics of the OPEC crude oil markets. Notable spikes in volatility coincide with major global events and financial disruptions, including the 1986 oil price collapse, the Gulf War in 1990, the 2008 global financial crisis, the 2014 oil price downturn (Baumeister and Kilian, 2016b; Prest, 2018), the COVID-19 pandemic in 2020 (Bourghelle et al., 2021; Le et al., 2023), and the Russia–Ukraine conflict (Zhang et al., 2023). These episodes, which triggered pronounced fluctuations in oil prices, are consistent with the historical patterns of oil shocks documented by Hamilton (2011).
@@
Figure 1: Panel (a) displays the monthly log returns and panel (b) shows the squared monthly log returns of oil prices from January 1983 to December 2022 for six selected OPEC countries: Algeria, Iran, Libya, Nigeria, Saudi Arabia, and the United Arab Emirates.
@@
Table LABEL:tab:stats reports the descriptive statistics for the monthly log returns of oil prices across the six major OPEC producers. All series exhibit positive means and medians, reflecting general upward trends but with varying dispersion levels, suggesting heterogeneous volatility. Negative skewness and high kurtosis are present in all cases, indicating asymmetric distributions with fat tails and substantial tail risk. These features—non-normality, asymmetry, and excess kurtosis—justify the use of heavy-tailed distributions in subsequent modeling.
@@
Figure LABEL:fig:hist complements these findings by illustrating empirical distributions. While returns are broadly centered around zero, the dispersion and shape vary markedly. Libya stands out with the widest and most asymmetric distribution, followed by Nigeria, with a moderate spread. In contrast, Saudi Arabia and the UAE show tighter and more concentrated return profiles. All histograms deviate from normality, as confirmed by the Jarque and Bera (1980) tests in Table LABEL:tab:stats. These stylized facts are consistent with the well-documented empirical properties of financial returns, such as time-varying volatility and leptokurtosis (Bera and Higgins, 1993).

### v2

# Forecasting Oil Volatility through Network Models with GARCH-Informed Correlation Weights

@@
Forecasting Oil Volatility through Network Models with GARCH-Informed Correlation Weights
@@
   Kahina Mehidi Department of Economics, University of Bejaia, Algeria. Email: kahina.mehidi@univ-bejaia.dz; ORCID: 0009-0001-3953-5155    Khelifa Mazouz Department of Economics, University of Bejaia, Algeria. Email: khelifa.mazouz@univ-bejaia.dz;
Cardiff Business School, Cardiff University, United Kingdom. Email: MazouzK@cardiff.ac.uk; ORCID: 0000-0001-6711-1715
   Philipp Otto School of Mathematics and Statistics, University of Glasgow, United Kingdom.
Email: Philipp.Otto@glasgow.ac.uk; ORCID: 0000-0002-9796-6682
This study addresses the computational challenges of forecasting volatility in high-dimensional commodity markets. Building on the Network log-ARCH framework, we introduce a novel class of network topologies from GARCH-informed correlation weights, obtained from conditional covariance estimates of multivariate GARCH models, rather than relying on the heuristic distance measures commonly used in clustering methods. We evaluate the proposed models forecasting performance through a rolling-window exercise using a panel of OPEC members crude oil prices. The results identify network volatility models incorporating these new GARCH-informed weights as the statistically superior specifications. Remarkably, the proposed framework matches standard DCC-GARCH predictive accuracy while delivering up to 62,000-fold computational gains. By explicitly modeling contemporaneous spillovers through interpretable spatial ARCH-like lags estimated via GMM, the proposed approach offers an optimal trade-off between parsimony, interpretability, and performance. The findings establish GARCH-informed network models as robust, scalable alternatives for systemic risk measurement and volatility forecasting in interconnected financial markets.
@@ 1 Introduction
Volatility forecasting poses a fundamental challenge across commodity markets, with petroleum markets showing exceptional complexity. The crude oil market ranks among the most volatile commodity markets (Regnier, 2007), and one of the most challenging assets to forecast. This volatility calls for innovative forecasting strategies and a deep understanding of price fluctuation mechanisms, which are crucial for practitioners and policymakers managing significant economic risks. The Organization of the Petroleum Exporting Countries (OPEC), one of the most influential organizations in global oil markets, plays a central role in market stabilization by coordinating production and pricing policies among member nations. Acting as a cartel, OPEC restricts oil supply to support price levels (Kilian, 2009; Al Rousan et al., 2018). Despite this coordination, significant heterogeneity persists across member countries regarding production behavior and pricing dynamics, implying complex interdependencies within the cartel structure. As a result, volatility transmission within OPEC may be highly non-trivial. Understanding these interdependencies is therefore critical for accurate volatility forecasting in global oil markets.
One promising approach to improving forecasts in interconnected markets is through network frameworks. Financial markets can be effectively represented as networks where asset returns serve as nodes and edges reflect the degree of similarity or connection between them (e.g. Barigozzi and Hallin, 2016; Betancourt et al., 2019; Demiris et al., 2013; Diebold and Yilmaz, 2014; Vinciotti et al., 2019; Zhou et al., 2023). Recent studies have increasingly adopted these network-based methodologies to explore volatility spillovers and market interconnectedness (e.g. Al Rousan et al., 2018; Ji et al., 2018; Baruník and Křehlík, 2018; Barigozzi and Brownlees, 2019; Bykhovskaya, 2023; Dai et al., 2023; Uddin et al., 2023; Barigozzi et al., 2025). Many of these models build upon the connectedness framework of Diebold and Yilmaz (2009, 2012, 2014), which quantifies directional spillovers using forecast error variance decompositions. However, these approaches often impose structure through rolling-window VARs and do not directly model the volatility process or incorporate the endogenous network structure of conditional heteroskedasticity, thereby limiting their ability to capture the structural mechanisms governing volatility transmission in highly interconnected markets.
Mattera and Otto (2024) address this limitation by introducing the Network Log-ARCH framework, which embeds network structure directly into the volatility process through spatial econometric methods. Their model captures contemporaneous volatility spillovers, where shocks in one market instantly affect volatility in connected markets, while maintaining computational tractability through GMM estimation. Applying this framework to stock market forecasting, they demonstrate superior performance over independent univariate log-ARCH models, with recent evidence suggesting that network-informed models can lead to better forecasts in financial markets (e.g. Wu et al., 2022; Huang et al., 2023). However, their analysis relies primarily on unconditional proximity measures, such as Euclidean distance and simple Pearson correlation, alongside K-nearest neighbors, to construct network weights. This approach does not leverage the rich information embedded in the conditional variance-covariance structure of financial returns, leaving open the question of whether economically informed network topologies could further enhance forecasting accuracy. Moreover, while several studies analyze volatility connectedness across oil markets, few have embedded empirical network structures directly into the volatility generating process (e.g. Bauwens et al., 2006; Ashfaq et al., 2020; Wang and Li, 2021; Sánchez García and Cruz Rambaud, 2023; Ben Salem et al., 2024).
This study addresses these gaps by extending the Network Log-ARCH framework in two key directions. First, we construct network weight matrices derived from the conditional correlation structures of standard MGARCH models, specifically CCC-GARCH, DCC-GARCH, and GO-GARCH. Unlike unconditional distance measures, these GARCH-informed networks encode the correlation structure implied by the underlying volatility dynamics of the system, capturing economic dependencies that purely geometric or unconditional statistical measures fail to detect. Importantly, we use time-averaged weights computed dynamically from the in-sample conditional correlations, providing stable network structures that reflect structural interdependencies without introducing look-ahead bias. Second, we apply this enhanced framework to OPEC oil markets, where production coordination and policy heterogeneity create unique spillover patterns distinct from equity markets. Building on the spatiotemporal ARCH modeling literature (Otto et al., 2024b), our approach enables modeling of structural volatility interdependence across a network of interconnected economies while directly inferring spillover dynamics from historical log return data.
Accordingly, this paper contributes to the literature in three main ways. First, it introduces GARCH-informed network weight matrices into the Network log-ARCH specification, derived from CCC, DCC, and GO-GARCH conditional correlations, allowing the model to exploit economically meaningful dependence structures. Second, it applies this framework to a panel of OPEC oil-exporting economies, providing new insights into the transmission of volatility among these interconnected member countries and demonstrating that network topology is a key determinant of forecasting performance. Third, it evaluates forecasting accuracy against standard multivariate GARCH specifications, demonstrating that GARCH-informed network models achieve comparable or superior predictive performance while drastically reducing computational cost, thereby offering a scalable alternative to traditional high-dimensional volatility models.
Our empirical findings confirm these advantages. Network Log-ARCH models incorporating CCC-, DCC-, and GO-GARCH-derived weight matrices yield lower Root Mean Squared Forecast Errors (RMSFE) and Mean Absolute Forecast Errors (MAFE) than Euclidean and correlation-based network alternatives, and perform competitively with, or better than, standard multivariate GARCH specifications. Predictive accuracy is evaluated using the Diebold and Mariano (2002) test and its heteroscedasticity-robust extension (Harvey et al., 2024). Furthermore, the best-performing models are rigorously identified through the Model Confidence Set (MCS) procedure of Hansen et al. (2011). Notably, the three GARCH-informed network specifications are jointly retained in the MCS at the 5% level and produce nearly identical point forecasts, while simpler network topologies (Euclidean, Pearson correlation, Piccolo distance) are excluded. These results highlight the central role of conditional correlation structures in shaping volatility dynamics across interconnected markets.
In the context of global energy markets, these models equip academic and professional energy economists, as well as policymakers, with a precise mechanism for tracking how price shocks propagate across major oil-exporting nations through contemporaneous spillovers. The network weights, particularly those derived from GARCH-informed conditional correlations, capture structural patterns in market interconnectedness. Crucially, this computational advantage transforms market surveillance. By shifting from retrospective to near real-time monitoring, it gives authorities a vital window to anticipate volatility shocks before they spread across the macroeconomic network. For economies heavily dependent on hydrocarbon revenues, such as OPEC member states, this enhanced volatility forecasting capability has direct implications for sovereign wealth management, macroeconomic planning, and the mitigation of risks stemming from systemic energy market shocks. Furthermore, these models provide practitioners and investors with highly efficient, advanced tools that drastically accelerate the forecasting process in network settings, supporting timely and informed financial decision-making under uncertainty.
@@
The remainder of this paper is organized as follows. Section 2 reviews the relevant literature on oil price volatility, forecasting, and network-based modeling. Section 3 describes the dataset and presents the key descriptive statistics. Section 4 provides an overview of multivariate volatility models. Section 5 outlines the methodological framework for the network-based log-ARCH models and the construction of the network weight matrices. Section 6 presents the empirical findings, including the estimation results, network structures highlighting interconnectedness and spillover effects, and the comparative forecasting performance of the proposed models; it concludes by outlining the advantages of using the proposed network volatility models.
@@ 2 Literature Review
Since its establishment in Baghdad in 1960, OPEC has aimed to coordinate and harmonize the petroleum policies of its member countries to ensure stable oil prices and adequate returns on investment (OPEC Statute, ). Over the decades, the organization has been shaped by a series of major historical and geopolitical events (OPEC Brief History, ), many of which have had far-reaching effects on the global economy (Jo, 2014). These events have exposed the persistent susceptibility of oil-exporting countries to external shocks, highlighting the complexity of managing coordinated policies in the face of heterogeneous economic responses (Liu et al., 2019). Geopolitical risks, particularly in oil-producing regions such as the Middle East and North Africa, have consistently driven oil price volatility (Mignon and Saadaoui, 2024). For instance, regional tensions and conflicts, along with global crises—such as the Chinese economic slowdown (Hamilton, 2014b)—have led to sharp fluctuations in oil prices. More recently, the COVID-19 pandemic amplified this volatility by triggering both supply- and demand-side shocks (Bourghelle et al., 2021; Le et al., 2023). Such compounded disruptions underscore the importance of understanding how volatility transmits across oil-exporting economies, particularly within policy-coordinated networks such as OPEC, where shocks may propagate unevenly across members.
@@
Oil price volatility remains a central concern in economic and financial forecasting, given its influence on key macroeconomic indicators, such as investment, durable goods consumption, and aggregate output. Uncertainty generated by oil price fluctuations can lead to reduced economic activity, particularly in resource-rich exporting nations (Finn, 2000; Kilian, 2008; Elder and Serletis, 2010; Vo and Le, 2021). These countries, as emphasized by Cappelli et al. (2023), are disproportionately affected due to their strong fiscal reliance on oil revenues. Despite decades of research, accurately forecasting oil price movements remains a formidable task due to complex interactions among market agents (Baumeister and Kilian, 2016).
@@
Conventional econometric models often fail to fully capture the complexity of oil price volatility, particularly its dynamic and interconnected nature within coordinated economic blocs. In response, Engle (1982) and Bollerslev (1986) developed the ARCH and GARCH models, which model volatility clustering in financial time series. These nonlinear frameworks remain foundational in financial econometrics, especially through multivariate extensions such as CCC-GARCH (Bollerslev, 1990), DCC-GARCH (Engle, 2002), and GO-GARCH (Van der Weide, 2002), which accommodate dynamic correlations in the return series. More recently, attention has shifted to models that incorporate spatial and network-based dependencies to account for cross-country volatility transmission. Spatiotemporal ARCH (spARCH) models incorporate contemporaneous spatial and temporal dependence to better capture volatility interlinkages (Otto et al., 2018), while spatial and spatiotemporal log-ARCH models (Sato and Matsuda, 2021; Otto et al., 2024b) and network log-ARCH models (Mattera and Otto, 2024) extend this framework to account for multiplicative effects. A comprehensive synthesis of these advanced approaches is presented in Otto et al. (2024a).
Empirical studies have shown strong interdependencies between oil prices and financial markets (Awartani and Maghyereh, 2013; Guhathakurta et al., 2020; Guan et al., 2024). For instance, Nguyen et al. (2020) document the short-run global repercussions of oil price shocks, while Creti et al. (2014) observe heightened oil–stock market co-movements in oil-exporting countries. However, existing studies often rely on models that assume static or exogenous relationships, limiting their ability to reflect the endogenous and evolving nature of the interdependencies among oil-exporting countries. In response, recent advances—such as time-series clustering methods (Maharaj et al., 2019) embedded into Dynamic Spatiotemporal ARCH models (Otto et al., 2024b)—offer promising tools for inferring network structures directly from return data. This opens new avenues for modeling volatility transmission in a way that reflects both temporal dynamics and cross-sectional spillovers, which is particularly relevant for OPEC economies, where coordinated production masks heterogeneous national responses. By integrating clustering-derived networks with advanced volatility models, our approach captures the latent and evolving interdependencies that shape volatility spillovers in oil-exporting nations. These methods move beyond traditional frameworks by eliminating the need for external variables, allowing for real-time analysis rooted in the structure of the return data itself. This offers both improved forecast accuracy and deeper insights into the underlying transmission channels relevant for policy coordination and risk management.
@@ 3 Data, Descriptive Statistics and Preliminary Analyses
The empirical analysis in this study is based on monthly crude oil price data for six key OPEC member countries: Algeria (Saharan Blend), Iran (Heavy), Libya (Es Sider), Nigeria (Bonny Light), Saudi Arabia (Arab Light), and the United Arab Emirates (Murban). These countries were selected based on the availability of a complete and consistent historical time series, which offered the same number of monthly observations across the entire sample. Crude oil prices were retrieved from the official OPEC database and subsequently converted into monthly logarithmic returns. The dataset spans the period from January 1983 to December 2024.
@@
The logarithmic returns and their squared values are presented in Figure 1, which illustrates the volatility dynamics of the OPEC crude oil markets over time. Notable spikes in volatility coincide with major global events and financial disruptions, including the 1986 oil price collapse, the Gulf War in 1990, the 2008 global financial crisis, the 2014 oil price downturn (Baumeister and Kilian, 2016; Prest, 2018), the COVID-19 pandemic in 2020 (Bourghelle et al., 2021; Le et al., 2023), and the Russia–Ukraine conflict (Zhang et al., 2023). These episodes, which triggered pronounced fluctuations in oil prices, are consistent with the historical patterns of oil shocks documented by Hamilton (2011a).
((a))
((b))
@@
Figure 1: Panel (a) displays the monthly log returns and panel (b) shows the squared monthly log returns of oil prices from January 1983 to December 2024 for six selected OPEC countries: Algeria, Iran, Libya, Nigeria, Saudi Arabia, and the United Arab Emirates.
@@
The descriptive statistics for the monthly log returns of prices across the six selected OPEC producers are presented in Table 1. All series exhibit positive means and medians, reflecting general upward trends but with varying dispersion levels, suggesting heterogeneous volatility. Negative skewness and excess kurtosis are present in all cases, indicating asymmetric distributions with fat tails and substantial tail risk. These features—non-normality, asymmetry, and leptokurtosis—justify the use of heavy-tailed distributions in subsequent modeling.
@@
The distributional plots in Appendix Figures B.1 and B.2 complement these findings by illustrating the empirical deviations. While returns are broadly centered around zero, the dispersion and shape vary markedly across countries. Libya exhibits the widest dispersion and the most extreme heavy tails, indicating a higher susceptibility to extreme shocks. Nigeria also displays a broad spread and the most pronounced asymmetry among the group. In contrast, Saudi Arabia and the UAE show tighter and more concentrated return profiles, though they still exhibit significant deviations from the Gaussian benchmark. All distributions decisively reject normality, as confirmed by the Jarque and Bera (1980) tests in Table 1. These stylized facts are consistent with the well-documented empirical properties of financial returns (Bera and Higgins, 1993).
Table 1: Descriptive Statistics of Monthly Oil Price Log Returns
Country Mean Median Std. Dev. Min. Max. Skew. Kurt. JB $p$-value
Algeria 0.0017 0.0066 0.0999 0.6952-0.6952 0.4957 0.683-0.683 8.04 <<2.2e-16***
Iran 0.0019 0.0053 0.1028 0.6475-0.6475 0.5536 0.605-0.605 6.82 <<2.2e-16***
Libya 0.0017 0.0071 0.1039 0.7700-0.7700 0.5215 0.798-0.798 9.89 <<2.2e-16***
Nigeria 0.0017 0.0088 0.1019 0.7654-0.7654 0.4708 0.838-0.838 9.67 <<2.2e-16***
Saudi Arabia 0.0018 0.0054 0.0975 0.6665-0.6665 0.5176 0.826-0.826 8.36 <<2.2e-16***
UAE 0.0017 0.0074 0.0893 0.4718-0.4718 0.4757 0.612-0.612 5.79 <<2.2e-16***
Notes: JB refers to the Jarque–Bera test for normality. *** indicates $p<0.001$. All series show non-normality with negative skewness and excess kurtosis.